OpenStudy (el_arrow):

find the sum of the series where x < 1

OpenStudy (el_arrow):

\[\sum_{n=0}^{\infty} x^n\]

OpenStudy (el_arrow):

@TheSmartOne could you help me out please

OpenStudy (el_arrow):

i know we can use any number greater than 1 but i dont know which one

TheSmartOne (thesmartone):

Hmm, I am not sure. Maybe @dan815 , @freckles , @mathmate , or @mathstudent55 can help you. :)

OpenStudy (el_arrow):

ok hopefully they can

OpenStudy (el_arrow):

hello mathmate

OpenStudy (mathmate):

Hello, This is a geometric series. @El_Arrow

OpenStudy (mathmate):

Are you familiar with geometric series, where the common ratio is x?

OpenStudy (el_arrow):

how did you know that right off the bat?

OpenStudy (el_arrow):

somewhat familiar we just started covering it today in class

OpenStudy (mathmate):

From one term to the next, you multiply by x, so that's the common ratio.

OpenStudy (el_arrow):

okay so its like a sequence

OpenStudy (mathmate):

Yes, it's a geometric sequence, but we're looking for the sum of the terms, so it is a series.

OpenStudy (mathmate):

If we sum to term N, then it becomes \(S=1+x+x^2+x^3+...,.x^N\) right?

OpenStudy (el_arrow):

yes

OpenStudy (mathmate):

Now do you know how to factorize: \(1-x^{N+1}\)

OpenStudy (el_arrow):

does it have to do something with dividing?

OpenStudy (mathmate):

Not in the factorizing part. Actually \(1-x^{N+1} = (1-x)(1+x+x^2+x^3+...x^N)\) If you're not sure, or have doubts, multiply it out to verify.

OpenStudy (mathmath333):

one method is this one \(\large \color{black}{\begin{align} y&=\sum_{n=0}^{\infty} x^n \hspace{.33em}\\~\\ y&=x^0+x^1+x^2+x^3+\cdots +\infty \hspace{.33em}\\~\\ y&=1+x^1+x^2+x^3+\cdots +\infty \hspace{.33em}\\~\\ y&=1+x\left(1+x^1+x^2+x^3+\cdots +\infty\right) \hspace{.33em}\\~\\ y&=1+xy \hspace{.33em}\\~\\ y-xy&=1 \hspace{.33em}\\~\\ y(1-x)&=1\hspace{.33em}\\~\\ y&=\dfrac{1}{1-x} \hspace{.33em}\\~\\ \sum_{n=0}^{\infty} x^n&=\dfrac{1}{1-x} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (el_arrow):

so the sequence will continue till it reaches x^n

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} y&=\sum_{n=0}^{\Large{\color{red}{\infty}}} x^n \hspace{.33em}\\~\\ \end{align}}\) the upper limit is \(\infty\) not \(n\) so it will go on forever

OpenStudy (mathmate):

Actually, if we divide both sides by (1-x), we get \((1-x^{N+1})/(1-x) = (1+x+x^2+x^3+...x^N)\) And the sum you are looking for can be found by putting N-> infinity. Since x<1, x^(N+1)=0, and you will get the result. The reason I take this approach is because sometimes you are required to sum to term N only, in which case you already for the formula for it.

OpenStudy (mathmate):

*already have the formula...

OpenStudy (el_arrow):

oh okay i see now

OpenStudy (el_arrow):

what if the n=1 would it still be the same formula

OpenStudy (mathmate):

Yes, exactly. Say, x=1/2 N=1 we have S=1+x=1+1/2=3/2 or \(S=(1-x^{N+1})/(1-x)\) \(=(1-(1/2)^2)/(1-1/2)\) \(=(1-1/4)/(1-1/2)\) \(=3/4*2/1\) \(=3/2\) as before.